### v1

# Quantum transistors for heat flux in and out of working substance parts: harmonic vs transmon and Kerr environs

Abstract
Quantum thermal transistors have been widely studied in the context of three-qubit systems, where each qubit interacts separately with a Markovian harmonic bath. Markovianity is an assumption that is imposed on a system if the environment loses its memory within short while, while non-Markovianity is a general feature, inherently present in a large fraction of realistic scenarios. Instead of Markovian environments, here we propose a transistor in which the interaction between the working substance and an environment comprising of an infinite chain of qutrits is based on periodic collisions. We refer to the device as a working-substance thermal transistor, since the model focuses on heat currents flowing in and out of each individual qubit of the working substance to and from different parts of the system and environment. We find that the transistor effect prevails in this apparatus and we depict how the amplification of heat currents depends on the temperature of the modulating environment, the system-environment coupling strength and the interaction time. We further show that there exists a non-zero amplification even if one of the environments, that is not the modulating one, is detached from the system. Additionally, the environment, being comprised of three-level systems, allows us to consider the effects of frail perturbations in the energy-spacings of the qutrit, leading to a non-linearity in the environment. We consider non-linearities that are either of transmon- or of Kerr-type. We find parameter ranges where there is a significant amplification for both transmon- and Kerr-type non-linearities in the environment. Finally, we detect the non-Markovianity induced in the system from a non-monotonic behavior of the amplification observed with respect to time, and quantify it using the distinguishability-based measure of non-Markovianity.
I Introduction
Advancements in science and technology have led to the development of modern equipment that requires miniaturized electrical circuits. These small-scale components exhibit thermal and quantum properties that impact performance of the device, making control over electrical and thermal fluctuations crucial. A transistor is a device that controls current flow in electric circuits Bardeen and Brattain (1948a); Millman and Halkias (1967, 1972); Boylestad and Nashelsky (2002). Electrical transistors are essential for modern technology, utilized in amplifiers, switches, and transformers. Since their invention by Bardeen and Brattain Bardeen and Brattain (1948a), extensive research has focused on improving and adapting transistor-based devices Millman and Halkias (1967, 1972); Boylestad and Nashelsky (2002); Bardeen and Brattain (1948b, 1949); Brinkman et al. (1997).
A quantum thermal transistor (QTT) Joulain et al. (2016) is the quantum analog of an electric transistor, used to control heat currents in miniaturized circuits. Quantum effects enable operations not possible in classical systems, offering advantages over traditional devices. Other quantum thermal machines like quantum batteries Alicki and Fannes (2013); Campaioli et al. (2018); Bhattacharjee and Dutta (2021), quantum refrigerators Palao et al. (2001); Feldmann and Kosloff (2003); Linden et al. (2010), quantum thermal diodes Yuan et al. (2021), and quantum rectifiers Werlang et al. (2014); Chioquetta et al. (2021); Poulsen et al. (2022) have been devised which exploit quantum mechanical properties of the system. Since the introduction of the quantum thermal transistor by Joulain et al. in 2016 Joulain et al. (2016), significant progress has been made, particularly in systems involving Markovian environments Guo et al. (2019); Naseem et al. (2020); Ghosh et al. (2021); Mandarino et al. (2021); Huangfu et al. (2021); Ruan et al. (2024); Yang and Tan (2020).
Current QTT models typically utilize a three-qubit working substance, each connected to a Markovian bath. Other models include qubit-qutrit systems Guo et al. (2018) and those employing periodic control, such as the Floquet quantum thermal transistor Gupt et al. (2022). There are extensive works which consider Coulomb-coupled conductors Yang et al. (2019); Zhang et al. (2018), polarons in the context of non-equilibrium systems Wang and Xu (2020), double cavity optomechanical systems Xiong et al. (2018), and three terminal normal-superconductors Tang et al. (2019) as models for QTTs. QTT performance, including environmental effects with and without noise, has been extensively studied in single and multi-transistor systems with shared baths. Ekanayake et al. (2023a); Liu et al. (2021); Ekanayake et al. (2023b); Mandarino (2022). The significance of quantum coherence in the context of QTT has also been studied, which enables to provide quantum advantage in the amplification of a QTT over its classical counterpart Su et al. (2021). Further studies on QTT focus on negative differential thermal conductance (NDTC) for high amplification rates and the role of strong coupling and quantum coherence in achieving NDTC Wang et al. (2018); Liu et al. (2019); Wu et al. (2024); Su et al. (2018). In cases requiring high amplification, transistors are often arranged in a Darlington configuration, where the output of one transistor connects to the input of another, greatly increasing the amplification factor. Models similar to these classical ones, but utilizing QTTs have also been rigorously explored. There are instances where such QTTs are placed in different configurations like the Darlington configuration Wijesekara et al. (2021, 2022) or a network of transistors, such as in Craven and Nitzan (2017), to attain many-fold advantage over scenarios that do not involve them. Experimental realization of QTTs using superconducting circuits Gubaydullin et al. (2022); Majland et al. (2020), materials like $VO_{2}$ Ordonez-Miranda et al. (2016); Prod’Homme et al. (2017), and implementation using magnetic thermal transistor Castelli et al. (2023) have been carried out.
Previous QTT studies utilize a global heat current approach, treating the system as three qubits where the focus is on the heat flowing from each bath into the entire working substance. The heat current is ”global” in a sense that the heat current corresponding to each of the baths flow globally into the entire working substance. This paper focuses on local heat currents, where we concentrate on the heat currents flowing “locally” into each qubit from different parts of the system and environment, in contrast to the global heat current approach. We have considered this approach since we know the characteristics of the sub-parts of the working substance experimentally, and it is essential to control and regulate the heat flow in the sub-parts of the system in order to preserve the material for actual implementations in circuits. Our concept of such local heat current is novel, and is hitherto unexplored in literature. There have been previous works which discuss the different types of heat currents in thermal equilibrium Hofer et al. (2017), where the system is connected to an Ohmic bath with a given spectral density. In Ref. Hewgill et al. (2021), for instance, it has been rigorously proven how local master equations are consistent with the laws of thermodynamics. As Markovian nature of a thermal environment, in general, arises under some specific assumptions, and is very unlikely to obtain in a practical scenario, it is important to study the effect of non-Markovianity on transistor action. Quantum thermal devices in the context of non-Markovian environments have been previously delved into Fischer and Breuer (2007); Bhattacharya et al. (2017); Bhattacharyya et al. (2025), which mainly focused on spin-environments. However, here we consider the system-environment interaction to be modeled by a collisional model comprising an infinite array of qutrits, which is experimentally feasible to implement and is fundamentally different from the Markovian environment scenario.
We provide a prototype of a QTT, which consists of three qubits interacting with each other and each qubit is connected separately to an environment. The interaction between the system and environment mimics the collisional model Bäumer et al. (2019); Scarani et al. (2002); Bruneau et al. (2014); Grimmer et al. (2016); Strasberg et al. (2017), which has been extensively studied in literature in the context of quantum batteries Sen and Sen (2023); Morrone et al. (2023), equilibrium and non-equilibrium dynamics Scarani et al. (2002); Bruneau et al. (2014); Grimmer et al. (2016); Seah et al. (2019); Cattaneo et al. (2021), strong coupling thermodynamics Strasberg (2019) and the impact of quantum coherence in thermodynamics Rodrigues et al. (2019); Hammam et al. (2021). In the model that we consider, each system qubit sequentially interacts with a qutrit of the environment for a short but fixed duration of time. The environment, which originally consists of infinite number three-level systems with equispaced energy levels, also allows us to consider the effects of frail perturbations in the energies of the environment-particles, leading to unequal energy spacings of each qutrit. Considering this model as the environment instead of the usual Markovian bosonic one, introduces an inherent non-Markovinity in the system. Using this model, we devise our transistor, which can control and regulate heat currents in the sub-parts of the working substance.
In this work, we scrutinize a class of transistors comprising a three-qubit working substance which interacts with the environment via the collisional model. We find significant amplification of the heat currents at two terminals of the transistor, being modulated by the heat current at the remaining terminal. The amplification factor depends on the temperature of the modulating environment, and system-environment coupling. Specifically, the amplification factors for the left and right qubits first increase and then decrease, as the modulating temperature rises. The amplification trend with coupling strength is complex: it decreases below a critical temperature but increases beyond it with stronger interactions. We also observe some amplification with fewer than three local environments. Additionally, in a linear system with equispaced energy levels, small fluctuations may occur due to external perturbations. Two examples of where such a situation naturally arises are transmons Koch et al. (2007); Josephson (1962); Ray et al. (2023a); Krause et al. (2022); Bhattacharyya et al. (2024) and Kerr-type environments Ray et al. (2023b). In our work, apart from considering non-Markovian environment, which constitutes the collisional model of qubits interacting sequentially with equispaced qutrits, we also consider the effects of inherent non-linearities in the environment such as in transmon and Kerr-types, in the amplification of a type of transistor. We find that amplification and its temperature dependence persist even with small non-linearities in the qutrit energy levels. Transmon and Kerr-type non-linearities enhance amplification in different temperature regimes. Below a critical temperature, Kerr non-linearities outperform linear qutrits, while above it, transmon qutrits are more effective. Furthermore, we find a non-monotonic behavior in the amplification factor of the three-qubit transistor setup with respect to time, and quantify the non-Markovianity that possibly induces such behavior.
The paper is organized as follows. In Sec. II, we explicitly describe the collisional model for the system and environment that we consider, along with defining the relevant Hamiltonians, the dynamical amplification factor, and the different types of heat-currents. We then analyze the effect of varying temperature of the modulating bath, and coupling constant between the bath and the working substance on the amplification considering linear baths in Sec. III. We consider both the cases where there are three (Sec. III.1) and less than three (Sec. III.2) environments. We then introduce non-linearity in the baths and analyze its effect on amplification in Sec. IV. In Sec. V, the non-Markovianity present in the system is quantified and analyzed. Finally the concluding remarks are presented in Sec. VI.
Figure 1: Schematic representation of the difference between the local and global approaches utilized to provide the definition of heat current used in previous models, referred to as BTTs versus the model studied in this paper, referred to as WTTs. The name, BTT essentially means that the transistor is conceptualized by looking at heat current flows in and out of each of the local baths to and from the entire system, as depicted by the red arrows in panel (a), and hence the transistor action is scrutinized by focusing on the “bath”. The name, WTT suggests that, the heat is exchanged between each (local) qubit with the different parts of the system and the environment, as shown by the blue arrows in panel (b), and therefore the focus is on the “working substance”. BTTs correspond to the usual QTT model previously studied in literature, while the class of transistor devices which we analyze in this paper are the WTTs.
II working-substance thermal transistor
An electrical transistor is a three terminal device which consists of a p-type (n-type) semiconductor sandwiched between two n-type (p-type) semiconductors to form a n-p-n (p-n-p) junction. The three terminals, corresponding to n-p-n, are often referred to as emitter, base and collector, and the currents flowing into each of these terminals are the emitter, base and collector currents, denoted by $I_{E},I_{B}$ and $I_{C}$ respectively. The function of an electrical transistor is to amplify the input base current to the output collector current, while operating in the common emitter mode. The figure of merit that quantifies the performance of a transistor is given by the dimensionless quantity, direct current (DC) amplification factor, $\beta_{DC}$, which is defined as
$\beta_{DC}=\frac{I_{C}}{I_{B}}.$
This figure of merit is useful, particularly in transistor circuits where all the currents are constant with respect to time. Meanwhile, if the currents are functions of time, the alternating current (AC) amplification factor, $\beta_{AC}$, is defined as
$\beta_{AC}=\frac{\Delta I_{C}}{\Delta I_{B}},$
where $\Delta I_{C}$ and $\Delta I_{B}$ are the small changes in $I_{C}$ and $I_{B}$ in the AC circuit of the transistor. The amplification factor refers to the change in collector current due to a small change in base current. Typically, the values of $\beta_{DC}$ or $\beta_{AC}$ is of the order of 200-300 and is an intrinsic property of an electrical transistor.
A QTT, on the other hand, is a thermal device which amplifies and modulates heat currents flowing into it, in analogy to an electrical transistor which amplifies electric currents. The quantum thermal transistor that we consider here consists of three two-level systems, which we refer to as the working substance. Each of these two-level systems is coupled separately to thermal environments. The interaction between the system and environment mimics that of the collisional model which we define in the succeeding subsection. We refer to the three two-level systems as $L,M,$ and $R$ which correspond to the left, middle, and right qubits, respectively. The energy differences between the two levels of the two-level systems are, respectively, given by $\hbar\omega_{L},\hbar\omega_{M}$ and $\hbar\omega_{R}$. The temperatures of the three thermal environments attached to the qubits are $T_{L},T_{M}$ and $T_{R}$ respectively. The middle environment is taken to be the modulating one which modulates the current in the left and right qubits analogous to its electrical counterpart. The dynamical amplification factor, $\alpha_{X}$ corresponding to the qubit-$X$, of a QTT is defined as
$\alpha_{X}=\frac{\partial J_{X}}{\partial J_{M}}=\frac{\frac{\partial J_{X}}{\partial T_{M}}}{\frac{\partial J_{M}}{\partial T_{M}}}$ (1)
where $X\in\{L,R\}$. The quantity, $J_{X}$, denotes the heat current flowing into the terminal $X$ of the working substance. A more elaborate discussion on heat currents, focusing on the type of heat current that we consider, is explicitly provided in Sec. II.3. The magnitude of amplification factor is indicative of how well the transistor performs, with the value of $|\alpha_{X}|$ being greater than $1$ indicating that the transistor effect exists.
II.1 Collisional model for system-environment interaction: linear and non-linear
The transistor models, that have been hitherto explored, fundamentally constitute local Markovian environments connected to each qubit of a QTT, where the notion of heat current corresponding to the individual heat flow from each bath was relevant. Here we use a different approach to design a type of transistor model, and provide a somewhat distinct definition of heat current, where we look into the heat current flowing into each of the qubits separately. This is discussed in detail in subsection II.3. Here we focus on the model of the environment that we consider. The working substance comprises three qubits, where each qubit is open to an environment. The environment constitutes an infinite number of three-level systems or qutrits. Each qutrit arrives one after the other, and interacts with a single qubit of the working substance for a fixed time, say $\delta t$. After each time interval, $\delta t$, a new qutrit pertaining to the environment arrives to interact with the qubit. Hence, the environment effectively resets after the finite interaction time, $\delta t$. Owing to the periodic collisions of the system with the environment, this model is often referred to in literature as the “collisional model” Bäumer et al. (2019); Scarani et al. (2002); Bruneau et al. (2014); Grimmer et al. (2016); Strasberg et al. (2017).
Initially we consider the energy difference between the ground and first excited states of each qutrit, that comprises the environment, to be equal to that between the first and second excited states of the qutrit. Since both the energy levels are equispaced, we refer to the type of environment comprising such qutrits to be a linear one. However in reality, there may be frail perturbations in the environment leading to small differences in the energy-level spacings. This introduces non-linearity in the environment due to a tweaking of the energy levels by a small amount. Let us say that the energy of the first excited level is increased by a small quantity. Then the energy gap between the ground state and the first excited state is greater than the difference between the first excited state and the second excited state. If the energy level differences decrease as we go to higher energy levels, then we get a transmon environment. Whereas if the energy level differences increase as we go from first to second excited states, a Kerr-type environment is generated. We refer to both these types of environments as non-linear ones.
II.2 Relevant Hamiltonians
In this subsection, we discuss the transistor model that we consider, including both the system and environment Hamiltonians, and the interactions between them. The Hamiltonian of the system comprising the three qubits is given by
$H_{sys}=-\sum_{i\in\{L,M,R\}}\frac{\hbar\omega_{i}}{2}\leavevmode\nobreak\ \sigma_{Z}^{i}-\sum_{\begin{subarray}{c}i,j\in\{L,M,R\}\\ i\neq j\end{subarray}}\frac{\hbar\omega_{ij}}{2}\leavevmode\nobreak\ \sigma_{Z}^{i}\leavevmode\nobreak\ \sigma_{Z}^{j},$ (2)
where $\sigma_{Z}$ refers to the Pauli-$z$ matrix for spin-$1/2$ particles. Here $L,M$ and $R$ refers to the left, middle and right qubits respectively. The first term represents the local Hamiltonian of each qubit, while the second term denotes the interaction between the qubits. The local Hamiltonian of a single qutrit of the environment is given by
$H_{env}=-\hbar\leavevmode\nobreak\ \Delta\sum_{i\in\{L,M,R\}}\sigma_{Z}^{(i),env},$ (3)
where the energy levels of each qutrit are equispaced, and $\Delta$ is the energy gap between any two successive energy levels. Here $\sigma_{Z}^{(i),env}$ is the Pauli-$z$ matrix for spin-$1$ particles. The index $i$ denotes the qutrit connected to $i^{\text{th}}$ qubit of the system. The Pauli-$x$ and Pauli-$z$ matrices for spin-$1$ are respectively given by
$\sigma_{x}^{env}=\frac{1}{\sqrt{2}}\begin{bmatrix}0&1&0\\ 1&0&1\\ 0&1&0\\ \end{bmatrix},\sigma_{z}^{env}=\begin{bmatrix}1&0&0\\ 0&0&0\\ 0&0&-1\\ \end{bmatrix}.$
In Sec. III, we consider the collisional model comprising of equispaced qutrits, which arrive one at a time to interact with the system qubit. We refer to this type of environment as linear because of the equal spacing between the energy levels of the three-level systems. Each of the three qubits interact locally with each qutrit of the environment via the qubit-qutrit interaction given by
$H_{qubit-env}=-\hbar\leavevmode\nobreak\ g\sum_{i\in\{L,M,R\}}\sigma_{X}^{(i),qubit}\leavevmode\nobreak\ \sigma_{X}^{(i),env}$ (4)
where $\sigma_{X}^{(i),qubit}$ is the Pauli-$x$ matrix for spin-$1/2$ particle, $\sigma_{X}^{(i),env}$ is the Pauli-$x$ matrix for spin-$1$ particle, and $g$ is the coupling constant between the system-qubit and the environment-qutrit. The quantities $\omega_{i},\omega_{ij},\Delta$ and $g_{k}$ are all in units of time-1, $\forall\leavevmode\nobreak\ i,j$.
For simplicity we consider $\omega_{L}=\omega_{M}=\omega_{R}=\omega_{ML}=\omega_{MR}=\Delta$ and $\omega_{LR}=0$. The disparity between the values of the interaction strengths $\omega_{ML}=\omega_{MR}$ and $\omega_{LR}$ introduces asymmetry in the system, which ensures transistor effect to be present.
II.3 Relevant heat currents in working-substance thermal transistors
The performance of a QTT depends on how efficiently it can amplify the heat currents in the terminals, $L$ and $R$, while being controlled by the flow of heat current at terminal $M$. The definition of heat current in the context of quantum transistors is given by the rate of change of energy of the working substance with respect to time, which focuses on the current flowing into and out of the individual baths. This scenario involves the consideration of local Markovian baths, where heat current is defined as
$\mathcal{J}_{X}=\text{Tr}(H_{sys}\mathcal{L}_{X}[\rho]),X\in\{L,M,R\},$
where $\rho$ is the density matrix of the working substance, and $\mathcal{L}_{X}[\rho]$ is the Lindblad operator corresponding to the terminal $X$, that appears in the Gorini-Kossakolski-Sudarshan-Lindblad (GKSL) Gorini et al. (1976); Lindblad (1976); Breuer and Petruccione (2002) master equation for the Markovian dynamics. We will refer to these transistors as bath thermal quantum transistors (BTTs) since in this case, one effectively considers the heat flowing in and out of a single bath into the entire working substance. This heat current is “global” in a sense that heat flows in and out of the entire system. We choose to focus on the heat current which flows in and out of each qubit of the working substance, to and from different parts of the system and environment, and refer to such type of transistors as working-substance thermal quantum transistors (WTTs) since the transistor action occurs due to the heat exchange between different parts of the system and environment, and a particular qubit. The local heat currents flowing in and out each $X^{\text{th}}$ terminal of the working substance is defined by
$J_{X}=\text{Tr}(\dot{\rho}_{X}H_{X}),$ (5)
where $X\in\{L,M,R\}$. The operator, $\dot{\rho}_{X}$, denotes the rate of change of the reduced density matrix of the $X^{\text{th}}$ terminal of the working substance. This type of heat current can be considered “local” in a sense that heat flows locally in and out of a particular subsystem of the working substance. The difference between the two approaches is pictorially depicted in Fig 1.
To find the amplification factor, we calculate the first order partial derivatives numerically using the five-point formula. According to the five-point formula, the first order derivative of a function $f(x)$ with respect to $x$ at point $x=x_{0}$ is given by
$f^{\prime}(x_{0})\\ =\frac{f(x_{0}-2h)-8f(x_{0}-h)+8f(x_{0}+h)-f(x_{0}+2h)}{12h},$ (6)
where $h$ is a small real number that is taken to be equal to $0.05$ in our calculations. The five-point formula introduces an error of $O(h^{4})$ which is equal to $6.25\times 10^{-6}$ in our case.
Figure 2: (a) Variation of $\partial J_{X}/\partial T_{M}$ with $T_{M}$, for $X\in\{L,M,R\}$, is depicted. Values of $\partial J_{M}/\partial T_{M}$ are much smaller than the other two, and becomes equal to zero at a particular value of $T_{M}$ which is equal to $6.65$ in units of $\tilde{T}$ when $g=4$. (b) Variation of $\partial{J_{M}}/\partial{T_{M}}$ with change in the coupling constant between the qubit and the environment is demonstrated. We see that, for every value of $g$, the quantity $\partial{J_{M}}/\partial{T_{M}}$ becomes zero at some value of $T_{M}=T_{M}^{critical}$, which makes the amplification factor undefined. This value of $T_{M}^{critical}$ increases as we increases the value of $g$. The black dotted line is for $\partial{J_{M}}/\partial{T_{M}}=0.$ The horizontal axis has dimension $\tilde{T}$ while the vertical axis when multiplied by $\hbar/\tilde{T}$ will have the dimension of ratio of rate of change of energy with temperature.
III WTT in presence of linear environment
In this section, we work with linear environment and consider the parameter space which allows the WIT to produce amplification. The temperature of the middle environment is taken to be in between that of the left and right environments. We consider the initial state of the system comprising the three qubits to be $|000\rangle\langle 000|$, where |$\ket{0}$ denotes the eigenstate of Pauli-$z$ with eigenvalue $+1$. The initial state of the environment $X$ is considered to be the thermal state, given by
$\rho_{env}^{X}=\frac{e^{\hbar\Delta\beta_{X}\sigma_{Z}^{(X),env}}}{\text{Tr}(e^{\hbar\Delta\beta_{X}\sigma_{Z}^{(X),env}})},$ (7)
where $\beta_{X}=1/(k_{B}T_{X})$ is the inverse temperature of the environment $X$ in units of energy-1, for $X\in\{L,M,R\}$.
We have defined time as $t\leavevmode\nobreak\ \tilde{t}$ where $t$ is a dimensionless quantity whose magnitude denotes the magnitude of time, and $\tilde{t}$ has the dimension of time with magnitude equal to $1$. Similarly, temperature is defined as $T\leavevmode\nobreak\ \tilde{T}$ where $T$ is dimensionless quantity denoting the magnitude of temperature, and $\tilde{T}$ has the dimension of temperature with magnitude equal to $1$. $\tilde{t}$ and $\tilde{T}$ are related by the relation
$\tilde{T}=\frac{\hbar}{k_{B}\leavevmode\nobreak\ \tilde{t}},$
where $k_{B}$ denotes the Boltzmann constant and $\hbar=h/2\pi$, with $h$ being the Planck’s constant.
The initial state of the system is considered to be $|000\rangle\langle 000|$, which is the ground state of the system Hamiltonian, $H_{sys}$. As we begin with the ground state of the system, it is expected that, during initial time evolution, heat will flow from the environment to the system. The sign convention adhered to in our work is that a positive heat current implies a flow of current from the system to the environment. This is what we numerically observe as well. The currents are initially negative at a small timescale $\sim 1\tilde{t}$. Further, since our system is not in a steady state, the relation of $J_{L}+J_{R}+J_{M}=0$ is not satisfied, and hence $\alpha_{R}+\alpha_{L}+1=0$ is not valid here.
The parameter range in which we perform the analyses is $T_{L}=4\tilde{T},T_{R}=10\tilde{T}$, $\Delta=3\tilde{t}^{-1},g=4\tilde{t}^{-1}$, the time of interaction of the system-qubit with the environment-qutrit is $\delta t=0.5\tilde{t}$, and the time at which amplification is measured is $t=1\tilde{t}$, unless specified. In Fig. 2, we demonstrate the behavior the derivatives of heat currents with respect to temperature, while operating within this particular parameter regime. Fig. 2-(a) shows that, $\partial J_{M}/\partial T_{M}$, the temperature derivative of the local current of middle qubit, is smaller than that of those corresponding to other qubits, within a temperature range, $4\tilde{T}$ to $10\tilde{T}$.
This enables us to obtain amplification within a sufficiently large temperature range for the middle environment. It is to be noted that we have chosen our parameter regime where the coupling strength between the environment and the working substance is comparable to the coupling strength between the sub-parts of the working substance. We see in panel Fig. 2-(a) that since $\partial J_{M}/\partial T_{M}$ is much smaller than the other two current derivatives in this parameter space, we get a good amplification. We also note that there exists a parameter region where $\partial J_{M}/\partial T_{M}=0$. Here the amplification is undefined. In panel (b) of Fig. 2, we have shown the variation of $\partial J_{M}/\partial T_{M}$ with $T_{M}$ for different values of system-environment coupling $g$.
In the succeeding subsection, we study the amplification of the heat currents flowing into the the left and right qubits in presence of local environments connected to each of the qubits separately. In subsection III.2, we analyze the thermal transistor effects with less than three local environments connected to three system qubits.
III.1 Amplification with three local environments
In this section, we consider the scenario where the system comprises three qubits, and each qubit is connected with a local environment that pertains to the collisional model composed of periodically interacting qutrits.
III.1.1 Dependence of amplification on the system-environment interaction
The performance of a WTT depends on how efficiently it can amplify the heat currents at the left and right terminals, while being modulated by the temperature of the middle terminal. Here we analyze how the amplification at the left and right terminals vary with temperature of the middle qubit by perturbatively varying the interaction strength between the working substance and environment.
Figure 3 illustrates the dynamical amplification factor at the left terminal, $\alpha_{L}$, as a function of the middle bath’s temperature, $T_{M}$, for different perturbative variations in the system-environment interaction strength, $g$. In this regard, we find that for a given interaction strength, the magnitude of amplification first increases until a certain value of $T_{M}$. Let us refer to this value as $T_{M}^{critical}$. At this particular temperature, $T_{M}^{critical}$, the dynamical amplification factor $\alpha_{L}$, diverges. The divergence occurs as a result of $\partial{J_{M}}/\partial{T_{M}}$ approaching zero, indicating a point of vanishing sensitivity of the current to changes in the temperature of the middle bath. This feature is depicted in Fig. 2-(b), where the derivative of the current with respect to $T_{M}$ is shown to diminish around $T_{M}^{critical}$. It is also observed that, at this point of discontinuity, the dynamical amplification factor, $\alpha_{L}$, switches its sign from negative to positive. Towards the right of $T_{M}^{critical}$, i.e. where $T_{M}>T_{M}^{critical}$, the magnitude of amplification decreases and saturates to a value which is greater than one. This illustrates that transistor effect is also present for relatively higher values of $T_{M}$. The value of $T_{M}^{critical}$ increases with the increase of system environment interaction strength. This feature can also be understood from Fig. 2-(b), where the value of $T_{M}$ for which the $\partial{J_{M}}/\partial{T_{M}}$ curve cuts the abscissa increases with increasing interaction strength. Furthermore, we show that, in the region $T_{M}<T_{M}^{critical}$, a lower interaction strength results in a higher magnitude of amplification being obtained for a particular value of $T_{M}$, whereas this trend is reversed in the $T_{M}>T_{M}^{critical}$ region, where a higher interaction strength results in higher amplification. The behavior of $\alpha_{R}$ with respect to $T_{M}$ is not shown because its behavior is identical to that of $\alpha_{L}$, and this occurs due to the symmetry between the left and right terminals of the setup.
Figure 3: Variation of the dynamical amplification factor, $\alpha_{L}$, along the vertical axis, versus the temperature of the middle bath, $T_{M}$, along the horizontal axis, for different values of system-bath interaction strength, $g$. We observe that there is a discontinuity in the value of $\alpha_{L}$ with respect to $T_{M}$, and an increase in the interaction strength increases and decreases the amplification to the right and left of the discontinuity respectively. The black dotted line signifies zero amplification. The quantity, $\alpha_{L}$, is dimensionless, while $T_{M}$ is in units of $\tilde{T}$.
III.1.2 Variation of amplification with time
In this section, we analyze the variation of dynamical amplification factor of a WTT as a function of time for a given value of system-environment coupling and a fixed temperature of the environment connected to the middle qubit. We find that the variation of the dynamical amplification factor with time becomes periodic after the occurrence of a certain number of interactions between the working substance and the environment. After a finite number of interactions, we find that the amplification shoots up and falls down periodically, as shown in Fig. 4. The curve obtained by joining the maxima of each jump, corresponding to a single round of system-environ interaction, is monotonically decreasing, which depicts Markovian behavior. However a non-monotonicity is observed in the actual behavior of amplification versus time, which owes its origin to the non-Markovianity induced in the system during the periodic interaction of the bath with the working substance. This non-monotonicity in the behavior of the amplification is because of the back-flow of information from the bath to the working substance, which occurs due to the non-Markovianity induced in the system during the periodic interaction of the bath with the working substance.
The jumps occur periodically with periodicity $0.5\tilde{t}$, which is the time of interaction between each qutrit of the environment and the working substance. As a qutrit pertaining to the environment arrives to interact with a system-qubit, the amplification remains small for a short time initially and then rises abruptly to a higher value. The maximum value corresponding to each jump is reached when the qutrit has stopped interacting with a qubit of the WTT. Following this, the amplification drops as the environment-qutrit moves away from the system-qubit. Then another qutrit arrives to interact with the qubit and the entire process is repeated. Here it is interesting to note that, in the time range in which we work, the number of peaks in the amplification versus time curve, as given in Fig. 4, is sufficient large. It basically implies a wide working region of the WTT in that particular range of time.
Figure 4: Variation of the dynamical amplification factor, $\alpha_{X}$, with time, $t$, at $T_{M}=10$. Here $X\in L,R$ where $L$ and $R$ denote left and right qubits respectively. We find that the variation becomes periodic after certain number of interactions have happened between the working substance and the environment. The dotted black line signifies amplification equal to zero. The vertical axis is dimensionless and the horizontal axis is in units of $\tilde{t}$.
III.2 Amplification with less than three local environments
In this section, we consider the scenario where the system comprises three or less qubits, and the number of environments interacting with the system are less than three.
III.2.1 WTTs with three qubits and two environments
To understand the role of the three environments further, we consider the case when we remove one of these environments. We have considered the middle environment, which is the modulating one, to be always present, and we remove either of the left or right environments. The temperatures of the three environments considered for the analysis in this subsection are given by $T_{L}=4\tilde{T},T_{M}=8\tilde{T},T_{R}=10\tilde{T}$.
[…]
Therefore, when the environment connected to the right qubit is removed, the value of $\alpha_{L}$ follows the same qualitative trend as that of $\alpha_{R}$ when the left bath is removed. But the amplification values are slightly higher in the first case of removing the right bath. The global maxima is reached at the same value of time but is higher and reaches a value of 13.84. The disparity between the two cases is because of the two baths having unequal temperatures which affect the amplification in each of the cases.
Figure 5: Variation of dynamical amplification factor with time when one of the environments is removed. The solid lines are for the system when the right environment is removed while the dotted lines are for the system when the left environment is removed . The inset shows the variation of amplification in the highlighted portion of the main graph. The vertical axis is dimensionless and the horizontal axis is in units of $\tilde{t}$.
III.2.2 Amplification with two qubits and two environments
In this subsection, we consider the case where the system comprises two qubits, left ($L$) and right ($R$), which are connected to respective environments having temperatures $T_{L}$ and $T_{R}$ (in dimensions of $\tilde{T}$) respectively. In this situation, we consider the left bath to be the modulating bath and see how a change in temperature of the left bath affects the amplification of the current in the right qubit. The dynamical amplification factor, in this case, is defined as
$\alpha=\frac{\partial J_{R}}{\partial J_{L}}=\frac{\frac{\partial J_{R}}{\partial T_{M}}}{\frac{\partial J_{L}}{\partial T_{M}}},$
where $J_{L}$ and $J_{R}$ are the heat currents flowing into the left and right qubits respectively. In this case, we observe that when the energy difference between the ground state and first excited state is the same for left and right qubit, then there is no amplification. However, when we consider unequal energy differences, there is a finite amplification in a small range of $T_{L}$. The range of $T_{L}$ is from 0 to 2.488 when $\omega_{L}=1$ and $\omega_{R}=2$, $T_{R}=4,g=4,\Delta=5$ and $t=1$.
IV Response to non-linear environments in amplification of WTT
Figure 6: Variation of the dynamical amplification factor, $\alpha_{L}$, along the vertical axis, versus the temperature of the middle bath, $T_{M}$, along the horizontal axis, for different values of non-linearity, given by $\epsilon$. In panel (a), variation of amplification ($\alpha_{L}$) with temperature of the middle bath, in the linear case along with the introduction of Kerr type and transmon non-linearities is given. The curves are obtained by taking time $t=1\tilde{t}$. The variation of amplification ($\alpha_{L}$) with temperature of the middle bath as we increase the non-linearity in the Kerr and transmon type environments at time $t=1\tilde{t}$ are plotted in panels (b) and (c) respectively. The quantity, $\alpha_{L}$, is dimensionless, while $T_{M}$ is in units of $\tilde{T}$.
Equispaced energy levels in a three-level system is a highly specific scenario and imposes strict constraints on the system, while the presence of non-linearity in systems is more probable in nature. Therefore, it is necessary to investigate the impact of non-linear environments on transistor action. Furthermore, there may arise frail perturbations in the energy levels of the environment qutrit, which generate non-linearity in the environment. In this section we analyze the dynamical amplification factor of a WTT in presence of non-linear environment. This means that the spacings between the energy levels are no longer equal as mentioned in Sec. II.
The local Hamiltonian of the environment connected to each qubit is given by
$H_{env}=-\hbar\begin{bmatrix}\Delta&0&0\\ 0&\epsilon&0\\ 0&0&-\Delta\\ \end{bmatrix},$ (8)
where $\epsilon$ is indicative of the amount of non-linearity in the environment. The condition, $\epsilon<0$, corresponds to the three-level system being a transmon qutrit, whose energy gaps gradually decrease as we go to higher energy levels. The condition, $\epsilon>0$, on the other hand, corresponds to Kerr type non-linearity of the environment-qutrit, where the energy gap increases as we consider higher energy levels. The quantity, $|\epsilon|$, can be considered to the strength of non-linearity (for either transmon or Kerr) present in the environment.
The three panels of Fig. 6 depict the behavior of the dynamical amplification factor for the left qubit, $\alpha_{L}$, as a function of the temperature of the middle environment, $T_{M}$, for different values of the non-linearity strength, $|\epsilon|$. In panel (a), the amplification for the linear case, transmon, and Kerr type environments are given by the purple, black and orange curves respectively. There is a discontinuity for all these three cases occurring at $T_{M}=T_{M}^{critical}$. The amplification, in each of the cases, increase as we increase $T_{M}$ when $T_{M}<T_{M}^{critical}$. We further find that in the low temperature regime, i.e. to the left of the discontinuity, Kerr type non-linearity provides a better amplification than the linear case, while in high temperature regime, i.e. to the right of the discontinuity, transmon qutrits provide an advantage in terms of better amplification than the linear case. In Fig. 6-(b), the dynamical amplification factor is plotted for different values of strength of non-linearity, $|\epsilon|$, and for $\epsilon>0$, i.e. in the Kerr-type regime. Similar to panel (a), a discontinuity in the behavior of amplification is observed. At low $T_{M}$ limit, to the left of the discontinuity of the orange curve, corresponding to $\epsilon=0.05$, an increase in the value of non-linearity strength increases the amplification for a fixed $T_{M}$. However, to the right of the discontinuity corresponding to the value curve corresponding to $\epsilon=0.01$, the feature is just the opposite, i.e. as we increase the strength of non-linearity, the amplification also increases.
In Fig. 6-(c), the dynamical amplification factor is plotted for different values of $|\epsilon|$, corresponding to $\epsilon<0$, i.e. in the transmon regime. Similar to the previous scenarios, here also a discontinuity in the behavior of amplification is noted. To the left of the discontinuity of black curve corresponding to $\epsilon=-0.01$, an increase in the value of the non-linearity strength decreases the amplification with $T_{M}$. However, to the right of the discontinuity of the orange curve corresponding to $\epsilon=-0.05$, the behavior is exactly the opposite and increasing the strength of non-linearity causes an increase in amplification.
Figure 7: Variation of the dynamical amplification factor, $\alpha_{L}$, with time, $t$. The inset shows the minute differences between the amplification in the case of transmon, Kerr, and linear environments in the range of time highlighted in the main graph. The vertical axis is dimensionless and the horizontal axis is in units of $\tilde{t}$. The given graph is at $T_{M}=10$ in units of $\tilde{T}$.
We further consider how the variation of amplification with time gets affected when non-linearity is introduced in the system. For the transmon and Kerr type environments, we observe that the amplification as a function of time, follows a trend which is similar to that of the linear environment case. However, the change in amplification when we consider non-linear environments is very small ($\sim 10^{-2}$) at all points of time except at time $t=\tilde{t}$, where the amplification has significant differences, i.e. $\epsilon=-0.01$ gives an amplification of $38.98$, $\epsilon=0.01$ gives an amplification of $34.00$, meanwhile linear type of environment gives an amplification of $36.27$. The maximum amplification that one can attain using this setup in presence of Kerr type environments is slightly higher $(38.98)$ than that using transmon qutrits $(34.00)$, albeit an amplification close to the linear case is obtained in both the cases.
V Quantifying non-Markovianity
As previously mentioned, the system that we consider deals with the global unitary evolution of the system and environment, followed by discarding the environment. This is, in general, non-Markovian in nature since there are no assumptions of Markovianity. Markovinity is a special case that arises when the environment does not possess the memory of its past events, i.e. due to a memoryless environment. In our case, when the environment interacts with the system, the memory of the environment changes depending upon the interaction. So this dynamics is non-Markovian at this stage, since the environment evolves depending upon its previous memory. However, the environment returns to its initial state periodically, i.e. there is also a loss of memory of the environment, which may potentially lead to Markovian behavior. Hence there is a tussle between memory-dependence and memory-independence of the environment in the transistor model that we have chosen. In this section, we study the memory-effects of the environment on the system, and quantify it in terms of the non-Markovinity induced in the system.
Different measures have been defined to detect non-Markovinity of a system undergoing a certain dynamics. We use the Breuer-Laine-Piilo (BLP) measure Breuer et al. (2009) - a faithful measure of non-Markovinity - to quantify non-Markovinity in our system. The trace distance, $D(\rho_{1},\rho_{2})$, between two states, $\rho_{1}$ and $\rho_{2}$, is a measure of how distinguishable the two states are from each other. It is defined by
$D(\rho_{1},\rho_{2})=\frac{1}{2}\text{tr}|\rho_{1}-\rho_{2}|,$
where the notation, $|A|=\sqrt{A^{\dagger}A}$, for any operator, $A$. The rate of change of trace distance between two states, denoted by $\sigma(t,\rho_{1,2}(0))$, is given by the following equation
$\sigma(t,\rho_{1,2}(0))=\frac{d}{dt}D(\rho_{1}(t),\rho_{2}(t)),$
where $t$ denotes time, and $\rho_{1(2)}(0)$ is the initial state of the $\rho_{1}(\rho_{2})$. If we begin with two distinguishable states, in a quantum Markovian process, since there is loss of information from the system throughout the process, the two states will become less distinguishable with time. It implies that the value of the quantity, $\sigma$, will have negative or zero value throughout the duration of the process. But in the case of a non-Markovian quantum process, since there is back flow of information from the environment to the system, we expect that the states will become more distinguishable and hence in a non-Markovian process, $\sigma$ will be positive and non-zero in a certain interval or throughout the process. The BLP measure $\mathcal{N}(\Phi)$ of non-Markovianity of a quantum process $\Phi$ is defined as
[…]
where the integration is over all positive values of $\sigma$, and the maximization is over all possible input states of $\rho_{1}$ and $\rho_{2}$.
To draw a parallel with the trends in amplification of the heat currents associated with each qubit, we consider the measure of non-Markovinity for each individual qubit of the working substance. We plot the measure $\mathcal{N}(\Phi)$ for each qubit, along the vertical axis, versus time $t$, along the horizontal axis in Fig. 8. In order to do so, we optimize over the state-space comprising of single qubits. It is to note that the optimal state-pairs belong to the boundary of the state-space Wißmann et al. (2012). Further, since we work with qubits, the boundary of the state-space contains only pure states, and therefore it is sufficient to optimize over pure states only. We consider pure states of the form
$\rho=\frac{1}{2}(\mathbb{I}+\vec{a}\leavevmode\nobreak\ \cdot\leavevmode\nobreak\ \vec{\sigma}),$ (10)
where $\vec{a}$ is the Bloch vector with $|\vec{a}|=1$, and the three components of $\vec{\sigma}$ are the spin-$\frac{1}{2}$ Pauli matrices, i.e. $\sigma_{X},\sigma_{Y}$, and $\sigma_{Z}$. The maximization is thus performed over the components of the Bloch vector with the constraint that the norm of the vector is equal to unity.
Figure 8: Plot of the BLP measure, $\mathcal{N}(\Phi)$, for each qubit of the working substance versus time, $t$. Here the dark blue, light blue, and grey curves correspond to the amplifications for the left, middle and right qubits of working substance respectively. We find that the non-Markovian effect exists up to time $t=1.5\tilde{t}$, after which the system behaves as Markovian.
Noting that the system is given periodic kicks of Markovianity through successive collisions between the system and environment, it is difficult to ascertain apriori, whether non-Markovian dynamics will dominate over Markovian. In Fig. 8, we observe the BLP measure for each qubit of the working substance, as a function of time. We find that at shorter timescales, there is back flow of information from the environment, and system is non-Markovian, as indicated by non-zero value of $\mathcal{N}(\Phi)$. However, at longer timescales, viz. after time $1.5\tilde{t}$, the BLP measure $\mathcal{N}(\Phi)=0$, implying that the system has reached a Markovian stage. This happens as the system loses memory of non-Markovian interactions at larger times. However, this Markovian feature is not reflected in the amplification (refer to Fig. 7), since there are non-monotonic jumps in the amplification even after time $1.5\tilde{t}$. This is because certain quantities, like the amplification in our case, has greater memory than the measure of non-Markovinity that we use here. The trend in the behavior of $\mathcal{N}(\Phi)$ is qualitatively the same for the left, middle and right qubits, as depicted in Fig. 8.
VI Conclusion
Quantum thermal transistors with Markovian baths have been rigorously studied and worked upon. Non-Markovian systems, however, are often more plausible in realistic situations. In this work, we considered a type of transistor which interacts with the environment via a collisional model, thereby deviating from Markovianity. We found that such an apparatus, which we referred to as a working-substance thermal transistor, does manifest the transistor effect. Such a nomenclature owes its origin to the fact that we focused on the heat currents which flow in and out of each qubit of the working substance due to the heat flow from and to the different parts of the system and environment, instead of on the currents which flow in and out of a single bath. We depicted how the figure of merit of the transistor, i.e. the amplification of the heat currents, vary with the temperature of the modulating environment, system-environment coupling and the interaction time. We have also analyzed the variation of amplification when one of the baths, that is not the modulating one, is removed.
Furthermore, we studied the effects of frail non-linearities in the environment, which are either of transmon- or Kerr-type. We found a significant amplification in each of these cases, and showed that both the type of non-linearities prove to be beneficial in different temperature regimes. In all of these cases, we also analyzed the variation of amplification with time. We found a non-monotonicity in the behavior of amplification versus time, which potentially is a manifestation of the inherent non-Markovianity in the system. We also quantified the amount of non-Markovianity by utilizing the distinguishability-based measure of Breuer, Laine and Piilo. Specifically, we found that the quantity of interest, i.e. the amplification, retains a memory of the non-Markovian interactions for a longer time than what the BLP measure itself is able to detect.
Acknowledgements.
We acknowledge computations performed using Armadillo Sanderson and Curtin (2016, 2018) on the cluster computing facility of Harish-Chandra Research Institute (HRI), India. We acknowledge the use of QIClib – a modern C++ library for general purpose quantum information processing and quantum computing. DB thanks HRI for support during visits.